heat in terms of kcals not kJ
How much heat is required to raise the temperature of 11.1 g of water from -45°C to 175 °C? heat: kcal
Answer #1

Specific heat capacity of water is 4.2 J/g o​​​​​​C.

So, for 1 gram water, to rise temperature by 1oC it needs 4.2 J heat.

Hence, for 11.1 g water, to rise temperature by 1oC it will need (4.2×11.1)J = 46.62 J heat.

Now, raising 1oC needs 46.62 J heat ( for 11.1g)

Hence, raising (175-(-45)) o​​​​​​C = 220oC will need (46.62×220) J = 10256.4 J heat = 10.2564 kJ heat.

Now, 1 kJ = 0.239 kcal

Hence, 10.2564 kJ = (0.239×10.2564)= 2.45 kcal.

Hence, the heat required is = 2.45 kcal.